Please help me with these???? Nobody answered the first time/ now I have to post it again and I lose a question :(

Please help me with these???? Nobody answered the
Answer #1

2 H2O2 =====> O2 + 2 H2O

Question A

30 gm of H2O2 = 30 /34.01 = 0.8819 Moles

2 Equivalent of H2O2 will produce 1 equivalent of O2 gas

0.8819 Moles of H2O2 will produce 0.4409 Moles of O2 gas

Question B

100 gm of H2O2 = 100 /34.01 = 2.9399 Moles

2 Equivalent of H2O2 will produce 2 equivalent of H2O

100 gm of H2O2 (2.9399 Moles) will produce 2.9399 Moles (52.96 gm) of H2O

Question C

2 Equivalent of H2O will need 2 equivalent of H2O2

50 Moles of H2O needs 50 moles of H2O2 which is 50 x34.01 = 1700 gram

Question d

50 gm of H2O2 = 50 /34.01 =1.4699 Moles

2 Equivalent of H2O2 will produce 1 equivalent of O2 gas

1.4699 Moles of H2O2 will produce 0.7349 Moles of O2 gas

0.7349 Moles of O2 gas = 0.7349 x 32 =  23.518 gram is theoretical yield

Percentage yield = 15 x100 / 23.518 = 63.78 %